OpenStudy (anonymous):

Of the 112 students in the marching band,35 were in the drum section .What is the ratio of drummers to other musicican in the band?

OpenStudy (anonymous):

35 to 77

OpenStudy (anonymous):

35/77 =5/11

OpenStudy (anonymous):

35:77 = 5:11

OpenStudy (anonymous):

35:77

OpenStudy (anonymous):

But ya simplified is 5:11 or 5/11 Possible options: 5:11 5/11 35 to 77

OpenStudy (anonymous):

Yay.

OpenStudy (anonymous):

Explanation: There are 112 students in all, but 35 are in drum section. So, without drums, there is 77 people. 112 - 35 35/77 = 5/11 35:77 = 5:11 Or 35 to 77

OpenStudy (anonymous):

thank your guy